Database last updated on Tue, 27 Jan 2026 15:40:51 +0000 by localhost
Critical warnings from b-files' comments for days between 2025108 and 2025108 (0 sec, 2 rows)
| Type | Comment | count | priority | fname | location | serial | time | value | min | max |
| rl: | 2025 107 05 35 13 Motor #2: appears stuck with a reference sensor activated. | 2 | 1 | B10825.224 | TORON | #224 | 05:27:30 | | 1 | -1 |
| rl: | 2025 108 21 10 45 New value set for watermark SUPPLY.VOLTAGE.LOW | 12 | 3 | B10825.224 | TORON | #224 | 05:27:35 | | 1 | -1 |
Instrument constant changes for the week before 2025108 (1 sec, 0 rows)
SL burned out? (AP results when SL is ON) (0 sec, 6 rows)
HG burned out? (AP results when HG is ON) (0 sec, 1 rows)
Non-critical warnings from b-files' comments for days between 2025108 and 2025108 (0 sec, 2 rows)
| Type | Comment | count | priority | fname | location | serial | time | value | min | max |
| rl: | 2025 107 05 33 50 Reset Warm -> Tepid from 2025 107 05 33 45 until 2025 107 05 33 45 . Counts: 1,0 | 1 | 3 | B10825.224 | TORON | #224 | 05:27:27 | | 1 | -1 |
| ds: | DS intensity too high ( 212227/cy) quitting. | 1 | 99 | B10825.224 | TORON | #224 | 20:48:32 | | 1 | -1 |
Measurements and tests summary.
| _jday | _loc | _serial | _skc | ds | dt | ed | fm | fz | hg | rs | sl | uv | zs |
| 108 | TORONTO | #224 | ecsc,Next schedule will be ecsc (now ecsc ZS= 93 ZS= 93.013) | 35 | 2 | 2 | 7 | 7 | 17 | 2 | 6 | 28 | 66 |
(1 sec)
Number of days since last dome/window cleaning (0 sec, 2 rows)
| serial | days |
| 224 | more than a year |
Missing files (0 sec, 1 rows)
| Serial | last data | days ago |
| 224 | 2025109 | 0-1 |
Run/Stop results for last week. (0 sec, 1 rows)
| Serial | Total | Failed 0 | Failed dark | Failed 1 | Failed 2 | Failed 3 | Failed 4 | Failed 5 | Failed double |
| 224 | 12 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
Standard lamp intensity analysis (0 sec, 1 rows)
| Serial | min | max | filter | fixable | 30-day history |
| #224 | 2666550 | 2667260 | 0 | yes |